# Hallucination Detection in LLMs with Topological Divergence on Attention Graphs

### v2

@@ C.1 Data Generation & Annotation
For human annotation, we asked three team members with at least upper-intermediate English proficiency to independently annotate approximately 100 samples from each dataset. We selected samples where all annotators reached a consensus and considered these annotations the ground truth hallucination labels.
@@
    Rouge-L scoring: we computed Rouge-L scores (using the evaluate library) between the model’s response and the ground-truth answers.
@@
Detailed statistics for each dataset can be seen in Table 8. The number of samples in the datasets varies across models, as we tried to maintain a balance of hallucinated and grounded responses, ensure sample cleanness, and minimize mislabeling. The procedure outlined above selects a different number of objects in a sample depending on the quality of the model’s responses.
@@
Appendix D Other experiment results
D.1 Alternative attention map-based features for hallucination detection
Figure 8: $\Delta_{ij}$ values for $ij$-th heads, MS MARCO vs CoQA. Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A). The heads that segregate samples best are highlighted in pink. Model names for a row are on the left side.
Features MS MARCO CoQA
Mistral-7B
Standard topological 0.67 0.69
Sparsity ratio 0.66 0.7
Entropy 0.75 0.77
Wasserstein 0.77 0.73
Spectral norm 0.73 0.72
Attention to <s> 0.65 0.61
MTop-Div 0.86 0.98
LLaMA-2-7B
Naive topological 0.69 0.7
Sparsity ratio 0.49 0.61
Entropy 0.38 0.68
Wasserstein 0.73 0.6
Spectral norm 0.49 0.64
Attention to <s> 0.62 0.64
MTop-Div 0.75 0.96
In our preliminary experiments for developing an attention maps-based hallucination detector, we trained classifiers using topological features previously applied to other NLP tasks Kushnareva et al. (2021); Cherniavskii et al. (2022), as well as traditional attention map characteristics. As standard topological features, we considered barcode-based features, such as the sum of bar lengths in persistence diagrams, and naive topological features, including the average vertex degree in attention graphs. For traditional attention-based features, we used sparsity ratio, attention entropy, and spectral norm Kobayashi et al. (2020); Vig and Belinkov (2019); Ji et al. (2021). We also considered Wassersein distances between the persistent diagrams Edelsbrunner and Harer (2010) of a context and a response subgraphs as an alternative way to assess their similarity. Finally, we analyzed the average attention to <s> token as we discovered that hallucination-aware heads often attend to it when the hallucination is present (see 4).
To determine the most informative features for hallucination detection, we trained supervised classifiers on concatenated features from all layers and heads and compared them to classifiers trained only on MTop-Div values under the same conditions. The results are presented in Table 9.
While the classifier based on MTop-Div values significantly outperforms alternative approaches, computing these values across all layers and attention heads is highly computationally expensive. To address this, we developed TOHA — a more efficient alternative by aggregating MTop-Div values from only a subset of the “hallucination-aware” attention heads in the model.
D.2 Other metrics for the head selection procedure
We also investigated alternative attention-map-based scores — including entropy, spectral norm, and the Wasserstein distance between the persistent diagrams of prompts and responses — for selecting specialized attention heads.Following the same pipeline as in TOHA, we computed the average distances between hallucinated and grounded samples using alternative proximity metrics. The results, presented in Figure 8, reveal that the most segregating heads for MS MARCO do not generalize to the CoQA dataset. This suggests that our proposed MTop-Div metric is better suited for the task compared to existing solutions.
Appendix E Implementation details
@@
  • For methods that rely on multiple generations, we generated 20 samples per input, following recommendations from Manakul et al. (2024); Chen et al. (2024).
  • For SelfCheck-GPT, we used its NLI-based variant.
@@
  • For Haloscope, we reserved 150 samples for hyperparameter tuning and treated the rest as unlabeled data. For the CoQA dataset, we added several thousand extra generations during training, as in the original paper. Other datasets were used as-is, since their sizes are comparable with those in Haloscope’s experiments on the TruthfulQA dataset.
  • For TOHA’s head selection, we similarly used 150 annotated samples. The topological divergences were calculated using ripser++ library Zhang et al. (2020), MIT license.
In our experiments, we used 60/15/25 train/val/test split. All the obtained results were averaged over five runs. All experiments were carried out using NVidia L40.
Appendix F Use of scientific artifacts & AI assistants
@@
The original text of this paper was spell- and grammar-checked and slightly smoothed out using Grammarly.
Appendix G Potential risks

### v3

@@ E.1 Data Generation & Annotation
For human annotation, we asked three team members with at least upper-intermediate English proficiency to independently annotate approximately 100 samples from each dataset. The human annotation consistency metrics are provided in Table 14. We selected samples where all annotators reached a consensus and considered these annotations the ground truth hallucination labels.
Pair Accuracy Pearson-r
1 vs 2 0.783 0.6
1 vs 3 0.812 0.631
2 vs 3 0.826 0.653
Mean 0.807 0.628
Std 0.022 0.027
Table 14: Cross-annotator consistency metrics.
@@
    Rouge-L scoring: we computed Rouge-L scores (using the evaluate library, v0.4.6) between the model’s response and the ground-truth answers.
@@
Detailed statistics for each dataset can be seen in Table 15. The number of samples in the datasets varies across models, as we tried to maintain a balance of hallucinated and grounded responses, ensure sample cleanness, and minimize mislabeling. The procedure outlined above selects a different number of objects in a sample depending on the quality of the model’s responses. Additionally, we provide the statistics regarding the response length distribution for each of the datasets we gathered (Figure 9).
@@
Figure 9: The distributions of response lengths (in words).
Figure 10: $\Delta_{ij}$ values for ijij heads, MS MARCO and CoQA. Vertical axis corresponds to the difference on the dataset (B), horizontal to the difference on the dataset (A). The heads that separate samples best are highlighted in pink.
Appendix F Implementation details
@@
  • For SelfCheckGPT, we used its NLI-based version.
@@
  • The topological divergences were calculated using ripser library Tralie et al. (2018), MIT license.
  • For the RAGTruth dataset, the model settings were aligned with the original paper.
All experiments were carried out using NVidia L40.
Appendix G Use of scientific artifacts & AI assistants
@@
The original text of this paper was spell- and grammar-checked and slightly smoothed using Grammarly.
Appendix H Potential risks